Calculate the maximum wavelength shift of a photon that undergoes Compton scattering from a free electron.


1
Expert's answer
2021-03-25T19:28:31-0400

"\lambda'-\lambda=\frac{h}{mc}(1-cos \theta),"

"(\lambda'-\lambda)_{max}=\frac{h}{mc}=\frac{6.63\cdot 10^{-34}}{9.11\cdot 10^{-31}\cdot 3\cdot 10^8}=2.43\cdot 10^{-12}~m."